# Adapting Multi-modal Large Language Model to Concept Drift From Pre-training Onwards

### v2

@@ Abstract
Multi-modal Large Language Models (MLLMs) frequently face challenges from concept drift when dealing with real-world streaming data, wherein distributions change unpredictably. This mainly includes gradual drift due to long-tailed data and sudden drift from Out-Of-Distribution (OOD) data, both of which have increasingly drawn the attention of the research community. While these issues have been extensively studied in the individual domain of vision or language, their impacts on MLLMs in concept drift settings remain largely underexplored. In this paper, we reveal the susceptibility and vulnerability of Vision-Language (VL) models to significant biases arising from gradual drift and sudden drift, particularly in the pre-training. To effectively address these challenges, we propose a unified framework that extends concept drift theory to the multi-modal domain, enhancing the adaptability of the VL model to the distribution unpredictable changes. Additionally, a T-distribution based drift adapter is proposed to effectively mitigate the bias induced by the gradual drift, which also facilitates the model in distinguishing sudden distribution changes through explicit distribution modeling. Extensive experiments demonstrate our method enhances the efficiency and accuracy of image-text alignment in the pre-training of VL models, particularly in the concept drift scenario. Moreover, various downstream tasks exhibit significant improvements in our model’s ability to adapt to long-tailed open world. Furthermore, we create a set of multi-modal datasets called OpenMMlo, specifically tailored for the long-tailed open world settings, to validate our findings. To foster the development of the multi-modal community, we have made both OpenMMlo datasets and our code publicly available at: https://github.com/Anonymous0Knight/ConceptDriftMLLMs.
@@ 1 Introduction
The rapid expansion of data availability has created significant challenges for multi-modal large language models (MLLMs), particularly in addressing concept drift, which predominantly manifests as gradual drift and sudden drift [1]. Among them, tailed drift represents a classic illustration of gradual drift, emerging due to severe data imbalance, where the distributions of long-tail categories evolves because of their intrinsic sparsity and noise. Concurrently, sudden drift is mainly represented by OOD drift, as the model encounters new, previously unseen concepts, resulting in distributional shifts that disrupt its ability to generalize in an open-world context. While the issues of long-tailed recognition and concept drift in open-world settings have been extensively studied in visual models [2] and language models [3], their impact on MLLMs, particularly vision-language (VL) models, remains largely unexplored. In this work, we aim to bridge this gap by providing a systematic analysis of how tailed drift and OOD drift affect VL models during both pre-training and fine-tuning phases. Our findings highlight critical vulnerabilities of current VL models in adapting to these challenges, underscoring the need for novel strategies to enhance their robustness in dynamic, open-world environments.
@@
Pre-training: As illustrated in Figure 1(a), a comparison of the VL model trained on the balanced dataset ImageNet [4] and the imbalanced dataset ImageNet-LT [2] is conducted. Due to the implicit feature centers of each category, we approximate them by averaging unit image and text features obtained by samples on the test set. To assess the intra-class compactness, the cosine distance between the image feature center and the text feature center from the same category is calculated and expressed as degrees. It is evident that training on the imbalanced dataset leads to a higher degree, indicating worse intra-class compactness brought by the tailed drift. Besides, with the tail drift intensifies, it results in a deterioration of the image-text alignment performance in tailed categories. Beyond the deterioration in tailed categories, the tailed drift also affects the image-text alignment in head categories with abundant training samples, which means that it leads to an overall performance degradation, not just the tailed categories. From the perspective of inter-class separability, we measure the average cosine distance from an image feature center to text feature centers of different categories. Figure 1(a) depicts that the VL model trained on the imbalanced dataset has lower inter-class separability. What’s more, we utilize KNN to extract 100 image and text feature centers of OOD samples to verify the impact of OOD drift on the pre-training of the VL model. Compared to training on the balanced dataset, the VL model trained under an imbalanced scenario is harder to distinguish between ID samples and OOD samples from the open world due to their similar inter-class separability. The undistinguished OOD drift will bias the underlying distribution of the feature space in the VL model, further disturbing the image-text alignment in the pre-training.
@@
Figure 1: The impacts of tailed drift and OOD drift on the vision language model in the stages of pre-training and fine-tuning, respectively. (a) In terms of the pre-training, we visualize the alignment results on both a balanced dataset (denoted as BL) and an imbalanced dataset (as LT). The cosine metric is used to measure the distances between unit image and text features across various categories including OOD samples, which is expressed as degrees. A smaller degree indicates a higher level of similarity between the features. Thus, it provides a feature-level visualization of the intra-class compactness and inter-class separability in the vision language model. (b) In the context of fine-tuning, the mutual cosine distance between the centers of each category in the classifier is directly visualized to illustrate the feature space of the classifier, denoted as blue bars. Besides, the average cosine distance between each category center and OOD samples is calculated, which is represented as orange bars.
Fine-tuning: We explicitly leverage the weights of the embedding layer in the VL model to visualize its feature space. The average cosine distance between each category and others is calculated as exhibited in the blue bars of Figure 1(b). with the decreasing of the training samples, a smaller degree means a worse inter-class separability. It is verified that tail drift leads to a compression of the feature space for tailed categories with a limited number of training samples, while head categories with abundant samples dominate the overall feature space of the VL model. Moreover, the average cosine distance between each category and unit features extracted by OOD samples is applied to reveal the OOD separability as denoted as orange bars in Figure 1(b). Since head categories occupies most of the feature space, OOD samples are closer to the center of the head categories compared to the tail categories, implying that in the stage of fine-tuning, it is difficult for the VL model to distinguish between OOD samples under imbalanced scenarios.
To effectively address tailed drift and OOD drift within a unified framework, which often occur simultaneously, we encapsulate them using the concept drift theory. Therefore, summarizing the above challenges of vision language models in the long-tailed open world, it raises the important question:
@@
Therefore, we propose a concept drift-aware multi-modal large language model, mitigating the tail drift and OOD drift encountered in the long-tailed open world. Firstly, we introduce the concept drift theory to the multi-modal domain, which provides a more holistic perspective to explain tailed drift and OOD drift. Then, the T-distributed adapter is proposed to be embedded in the hyperspherical feature space. It aligns image-text features for contrastive learning in pre-training. The desirable light-tailed property of the proposed T-distributed spherical metric (Thp) prevents the compression of tailed categories and mitigates the crowding of feature space caused by tailed drift. Besides, in fine-tuning, the adapter projects features to decision boundary and detects OOD samples at feature-level based on the underlying distribution. The proposed T-hp distribution explicitly models the feature space with concrete feature centers, optimizing large inter-class margins and yielding more desirable hyperspherical embeddings. And a simple non-parametric KNN is adopted to distinguish the OOD sample based on the T-hp distribution. Finally, we construct a group of multi-modal long-tailed open datasets to support our claims.
@@
    We build a group of multi-modal datasets OpenMMlo under the long-tailed open world, by extending existing image-based long-tailed open datasets. It contains about 740k image-caption pairs with related category annotations. To support and encourage the community focused on multi-modal, we have made both the OpenMMlo and our code public.
@@ 2.1 Multi-modal Concept Drift Theory
Concept drift is a phenomenon in which the statistical properties of a target domain change over time in an arbitrary way [1]. Formally, given a set of examples denoted as the data stream $S_{0,t}=\{d_{0},...,d_{t}\}$, where $d_{i}=(X_{i},y_{i})$ is one data instance, $X_{i}$ and $y_{i}$ respectively denote the feature vector and the label, and $t$ represents the timestamp of the instance in the data stream. $S_{0,t}$ follows a certain distribution $F_{0,t}(X,y)$. The concept drift is formalized as: $\exists t:P_{t}(X,y)\neq P_{t+1}(X,y)$, where the joint probability $P_{t}(X,y)$ can be decomposed as $P_{t}(X,y)=P_{t}(X)\times P_{t}(y|X)$. Although the concept drift due to tailed and OOD data often co-occur, they are fundamentally distinct phenomena. The tailed concept drift foucus on the drift in $P_{t}(X)$, while $P_{t}(y|X)=P_{t+1}(y|X)$ remains unchanged. However, the OOD drift from the unknown categories triggers the drift of both $P_{t}(y|X)$ and $P_{t}(X)$, that $P_{t}(y|X)\neq P_{t+1}(y|X)$ and $P_{t}(X)\neq P_{t+1}(X)$. Therefore, concept drift theory provides a unified framework to harmonize the tailed shift and OOD shift that often occur together, enabling more robust and adaptive deep learning models.
In the context of multi-modal vision language models, we extend the concept drift theory from a single data stream to multiple data streams. Each data modal is associated with a distinct data stream. Thereby, the multi-modal concept drift framework can robustly handle the complex, heterogeneous data distributions inherent to vision language models. Therefore, we formally define multi-modal concept drift as follows:
@@
$\exists t:P_{t}(S_{i})\neq P_{t+1}(S_{i})$ (1)
Remark 2.1.
Figure 2: The workflow of our methodology, consisting of two stages: the pre-training of the vision-language model and the fine-tuning on downstream tasks. Within the data streaming, a drift adaptation window slides to detect changes in data distribution and subsequently update the model, in both pre-training and fine-tuning. In the pre-training, the T-distributed adapter aligns visual and textual feature space by image-text contrastive learning, with a large inter-class margin. Coupled with the language model loss, they drive the training of all modules. In the downstream task, the image encoder and the text decoder are frozen out of training, with a linear projector fusing image-text features. Additionally, a mixture of experts module is leveraged with the T-distributed adapter as the router, allowing it to effectively adapt tail drift and perform OOD drift detection based on the distribution.
@@ 2.2 T-distributed Adapter for Concept Drift
To adapt the vision language model to concept drift, it is essential to adapt the model to align with the evolving data distribution, which can be formally defined as
$\min_{f^{(t)},f^{(t+1)},...,f^{(t+\tau)}}\sum_{i=t}^{t+\tau}L(f^{(i)}(x^{(i)}),y^{(i)})$ (2)
where $f^{(t)}$ denotes the vision language model trained by the data stream $S_{t-k,t-1}$ from the drift adaption window with the size of $k$. And the model is driven by the target metric $L$ continuously to adapt the drift in a given time period $[t,t+\tau]$. Thus, one prevalent method for detecting and adapting to concept drift involves designing metrics based on data distributions that can effectively counteract the impacts of sudden and gradual changes within the time window [5, 6]. Building upon this thinking, we integrate directional statistics into distribution-based drift detection and adaptation, proposing a T-distributed adapter to alleviate it in the vision language model. Firstly, we provide an overview of directional statistics in the Appendix LABEL:appendix:ds. Then, we will introduce the T-distributed adapter.
Inspired by the T-SNE [7], we design a T-distribution based metric in hypersphere (T-hp), which follows the density:
$p_{X}(x^{(i)})\propto\frac{2}{\kappa(1-\mu^{T}x^{(i)})}$ (3)
@@
$N_{T}(\kappa,d)=\int_{\mathbb{S}^{d-1}}{\frac{2}{1-\kappa\mu^{T}x^{(i)}}}\mathrm{d}x=\frac{1}{\kappa}2^{\alpha+\beta-1}\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)}$ (4)
where $\alpha=\frac{d-1}{2}$, $\beta=\frac{d-3}{2}$, and $\Gamma(\cdot)$ represents the gamma function. Combined with Eq. LABEL:eq:2, the normalizer $N_{X}(d)$ of density $p_{X}(x;\mu)$ is:
$N_{X}(\kappa,d)=N_{T}(\kappa,d)\cdot A_{d-2}=\frac{2^{\alpha+\beta}\pi^{\beta}}{\kappa}\frac{\Gamma(\alpha)}{\Gamma(\alpha+\beta)}$ (5)
@@
$p(x^{(i)})=N_{X}(d)^{-1}\frac{2}{\kappa(1-\mu^{T}x^{(i)})},\quad x^{(i)}\sim\text{Thp}(\mu)$ (6)
The detailed derivation process is provided in Appendix LABEL:appendix:Tsp.
@@
In terms of adapting to tailed concept drift, the Thp metric with a large concentration exhibits a light-tailed property, wherein the probability density function exhibits a faster rate of decay as the values increase, relative to the vMF metric, as illustrated in Figure 3. The high kurtosis of Thp is characterized that it yields high confidence only when the feature vector is sufficiently close to the center of the category, thereby minimizing the influence of head category samples on the tail category centers. Formally, $L_{\text{thp}}(\mu,x^{(i)})=\frac{2}{\kappa(1-\mu^{T}x^{(i)})+\epsilon}$ denotes the T-distributed metric on hypersphere, where $\epsilon$ is a non-zero value setting to $1$ to avoid the denominator of 0, and $L_{\text{vmf}}(\mu,x)=\exp(\kappa\mu^{T}{x})$ represents the vMF metric. Given an unit feature vector $x_{\text{head}}$ from the head categories, the gradient of the metric $L$ over the tailed category center $\mu_{\text{tail}}$ is $\frac{\partial L(\mu_{\text{tail}},x_{\text{head}})}{\partial\mu_{\text{tail}}}$. Due to $\mu^{T}_{\text{tail}}x_{\text{head}}\in[-1,1]$, when $\kappa\geqslant 1$, it is readily obtain that $\frac{\partial L_{\text{thp}}}{\partial\mu_{\text{tail}}}<\frac{\partial L_{\text{vmf}}}{\partial\mu_{\text{tail}}}$. Consequently, the light-tailed Thp distribution effectively counteracts the squeezing of tail categories caused by an overwhelming number of head samples, thereby alleviating the bias induced by tailed concept drift.
Likewise, the Thp metric is directly applied to detect the OOD concept drift. A sample with a unit feature vector $x$ is deemed out-of-distribution if it lies at a relatively large distance from the in-distribution (ID) data in the eigenspace. Following [8], a simple non-parametric KNN is adopted to partition the data into two sets (ID vs. OOD), which does not impose any distributional assumption on the feature space. Here the distance is the Thp metric with respect to the k-th nearest neighbor.

### v3

@@ Abstract
Multi-modal Large Language Models (MLLMs) frequently face challenges from concept drift when dealing with real-world streaming data, wherein distributions change unpredictably. This mainly includes gradual drift due to long-tailed data and sudden drift from Out-Of-Distribution (OOD) data, both of which have increasingly drawn the attention of the research community. While these issues have been extensively studied in the individual domain of vision or language, their impacts on MLLMs in concept drift settings remain largely underexplored. In this paper, we reveal the susceptibility and vulnerability of Vision-Language (VL) models to significant biases arising from gradual drift and sudden drift, particularly in the pre-training. To effectively address these challenges, we propose a unified framework that extends concept drift theory to the multi-modal domain, enhancing the adaptability of the VL model to unpredictable distribution changes. Additionally, a T-distribution based drift adapter is proposed to effectively mitigate the bias induced by the gradual drift, which also facilitates the model in distinguishing sudden distribution changes through explicit distribution modeling. Extensive experiments demonstrate our method enhances the efficiency and accuracy of image-text alignment in the pre-training of VL models, particularly in the concept drift scenario. Moreover, various downstream tasks exhibit significant improvements in our model’s ability to adapt to the long-tailed open world. Furthermore, we create a set of multi-modal datasets called OpenMMlo, specifically tailored for the long-tailed open-world setting, to validate our findings. To foster the development of the multi-modal community, we have made both OpenMMlo datasets and our code publicly available at: https://github.com/XiaoyuYoung/ConceptDriftMLLMs.
@@ 1 Introduction
The rapid expansion of data availability has created significant challenges for multi-modal large language models (MLLMs), particularly in addressing concept drift, which predominantly manifests as gradual drift and sudden drift Lu et al. (2019). Among them, tailed drift represents a classic illustration of gradual drift, emerging due to severe data imbalance, where the distributions of long-tail categories evolve because of their intrinsic sparsity and noise. Concurrently, sudden drift is mainly represented by OOD drift, as the model encounters new, previously unseen concepts, resulting in distributional shifts that disrupt its ability to generalize in an open-world context. While the issues of long-tailed recognition and concept drift in open-world settings have been extensively studied in visual models Liu et al. (2022b) and language models Kandpal et al. (2023), their impact on MLLMs, particularly vision-language (VL) models, remains largely unexplored. In this work, we aim to bridge this gap by providing a systematic analysis of how tailed drift and OOD drift affect VL models during both pre-training and fine-tuning phases. Our findings highlight critical vulnerabilities of current VL models in adapting to these challenges, underscoring the need for novel strategies to enhance their robustness in dynamic, open-world environments.
@@
Pre-training: As illustrated in Figure 1(a), a comparison of the VL model trained on the balanced dataset ImageNet Russakovsky et al. (2015b) and the imbalanced dataset ImageNet-LT Liu et al. (2022b) is conducted. Due to the implicit feature centers of each category, we approximate them by averaging unit image and text features obtained by samples on the test set. To assess the intra-class compactness, the cosine distance between the image feature center and the text feature center from the same category is calculated and expressed as degrees. It is evident that training on the imbalanced dataset leads to a higher degree, indicating worse intra-class compactness brought by the tailed drift. Besides, with the tail drift intensifies, it results in a deterioration of the image-text alignment performance in tailed categories. Beyond the deterioration in tailed categories, the tailed drift also affects the image-text alignment in head categories with abundant training samples, which means that it leads to an overall performance degradation, not just the tailed categories. From the perspective of inter-class separability, we measure the average cosine distance from an image feature center to text feature centers of different categories. Figure 1(a) depicts that the VL model trained on the imbalanced dataset has lower inter-class separability. What’s more, we utilize KNN to extract 100 image and text feature centers of OOD samples to verify the impact of OOD drift on the pre-training of the VL model. Compared to training on the balanced dataset, the VL model trained under an imbalanced scenario is harder to distinguish between ID samples and OOD samples from the open world due to their similar inter-class separability. The undistinguished OOD drift will bias the underlying distribution of the feature space in the VL model, further disturbing the image-text alignment in the pre-training.
@@
Figure 1: The impacts of tailed drift and OOD drift on the vision language model in the stages of pre-training and fine-tuning, respectively. (a) In terms of the pre-training, we visualize the alignment results pre-trained on both a balanced dataset (denoted as BL) and an imbalanced dataset (as LT) without OOD samples, under the same balanced test set.The cosine metric is used to measure the distances between unit image and text features across various categories including OOD samples, which is expressed as degrees. A smaller degree indicates a higher level of similarity between the features. Thus, it provides a feature-level visualization of the intra-class compactness and inter-class separability in the vision language model. (b) In the context of fine-tuning in imbalance datasets, the mutual cosine distance between the centers of each category in the classifier is directly visualized to illustrate the feature space of the classifier, denoted as blue bars. Besides, the average cosine distance between each category center and OOD samples is calculated, which is represented as orange bars.
Fine-tuning: We explicitly leverage the weights of the embedding layer in the VL model to visualize its feature space. The average cosine distance between each category and others is calculated as exhibited in the blue bars of Figure 1(b). With the decreasing of the training samples, a smaller degree means a worse inter-class separability. It is verified that tail drift leads to a compression of the feature space for tailed categories with a limited number of training samples, while head categories with abundant samples dominate the overall feature space of the VL model. Moreover, the average cosine distance between each category and unit features extracted by OOD samples is applied to reveal the OOD separability as denoted as orange bars in Figure 1(b). Since head categories occupy most of the feature space, OOD samples are closer to the center of the head categories compared to the tail categories, implying that in the stage of fine-tuning, it is difficult for the VL model to distinguish between OOD samples under imbalanced scenarios.
To effectively address tailed drift and OOD drift within a unified framework, which often occurs simultaneously, we encapsulate them using the concept drift theory. Therefore, summarizing the above challenges of vision language models in the long-tailed open world, it raises the important question:
@@
Therefore, we propose a concept drift-aware multi-modal large language model, mitigating the tail drift and OOD drift encountered in the long-tailed open world. Firstly, we introduce the concept drift theory to the multi-modal domain, which provides a more holistic perspective to explain tailed drift and OOD drift. Then, the T-distributed adapter is proposed to be embedded in the hyperspherical feature space. It aligns image-text features for contrastive learning in pre-training. The desirable light-tailed property of the proposed T-distributed spherical metric (Thp) prevents the compression of tailed categories and mitigates the crowding of feature space caused by tailed drift. Besides, in fine-tuning, the adapter projects the features to the decision boundary and detects OOD samples at the feature level based on the underlying distribution. The proposed T-hp distribution explicitly models the feature space with concrete feature centers, optimizing large inter-class margins and yielding more desirable hyperspherical embeddings. And a simple non-parametric KNN is adopted to distinguish the OOD sample based on the T-hp distribution. Finally, we construct a group of multi-modal long-tailed open datasets to support our claims.
@@
    We build a group of multi-modal datasets OpenMMlo under the long-tailed open world by extending existing image-based long-tailed open datasets. It contains about 740k image-caption pairs with related category annotations. To support and encourage the community focused on multi-modal, we have made both the OpenMMlo and our code public.
@@ 2.1 Multi-modal Concept Drift Theory
Concept drift is a phenomenon in which the statistical properties of a target domain change over time in an arbitrary way Lu et al. (2019). Formally, given a set of examples denoted as the data stream $S_{0,t}=\{d_{0},...,d_{t}\}$, where $d_{i}=(X_{i},y_{i})$ is one data instance, $X_{i}$ and $y_{i}$ respectively denote the feature vector and the label, and $t$ represents the timestamp of the instance in the data stream. $S_{0,t}$ follows a certain distribution $F_{0,t}(X,y)$. The concept drift is formalized as: $\exists t:P_{t}(X,y)\neq P_{t+1}(X,y)$, where the joint probability $P_{t}(X,y)$ can be decomposed as $P_{t}(X,y)=P_{t}(X)\times P_{t}(y|X)$. Although the concept drift due to tailed and OOD data often co-occur, they are fundamentally distinct phenomena. The tailed concept drift foucus on the drift in $P_{t}(X)$, while $P_{t}(y|X)=P_{t+1}(y|X)$ remains unchanged. However, the OOD drift from the unknown categories triggers the drift of both $P_{t}(y|X)$ and $P_{t}(X)$, that $P_{t}(y|X)\neq P_{t+1}(y|X)$ and $P_{t}(X)\neq P_{t+1}(X)$. Therefore, concept drift theory provides a unified framework to harmonize the tailed shift and OOD shift that often occur together, enabling more robust and adaptive deep learning models.
In the context of multi-modal vision language models, we extend the concept drift theory from a single data stream to multiple data streams. Each modality is associated with a distinct data stream. Thereby, the multi-modal concept drift framework can robustly handle the complex, heterogeneous data distributions inherent to vision language models. Therefore, we formally define multi-modal concept drift as follows:
@@
$\exists t:P_{t}(S_{i})\neq P_{t+1}(S_{i}).$ (1)
Figure 2: The workflow of our methodology, which consists of two stages: the pre-training of the vision-language model and the fine-tuning on downstream tasks. Within the data streaming, a drift adaptation window slides to detect changes in data distribution and subsequently update the model, in both pre-training and fine-tuning. In the pre-training, the T-distributed adapter aligns visual and textual feature space by image-text contrastive learning, with a large inter-class margin. Coupled with the language model loss, they drive the training of all modules. In the downstream task, the image encoder and the text decoder are frozen out of training, with a linear projector fusing image-text features. Additionally, a mixture of expert modules is leveraged with the T-distributed adapter as the router, allowing it to effectively adapt tail drift and perform OOD drift detection based on the distribution.
@@ 2.2 T-distributed Adapter for Concept Drift
To adapt the vision language model to concept drift, it is essential to adapt the model to align with the evolving data distribution, which can be formally defined as:
$\min_{f^{(t)},f^{(t+1)},...,f^{(t+\tau)}}\sum_{i=t}^{t+\tau}L(f^{(i)}(x^{(i)}),y^{(i)}),$ (2)
where $f^{(t)}$ denotes the vision language model trained by the data stream $S_{t-k,t-1}$ from the drift adaption window with the size of $k$. And the model is driven by the target metric $L$ continuously to adapt the drift in a given time period $[t,t+\tau]$. Thus, one prevalent method for detecting and adapting to concept drift involves designing metrics based on data distributions that can effectively counteract the impacts of sudden and gradual changes within the time window Jiao et al. (2022); Yu et al. (2024). Building upon this thinking, we integrate directional statistics into distribution-based drift detection and adaptation, proposing a T-distributed adapter to alleviate it in the vision language model. Firstly, we provide an overview of directional statistics in the Appendix A.2.1. Then, we will introduce the T-distributed adapter.
Inspired by the T-SNE Van der Maaten & Hinton (2008), we design a T-distribution based metric in hypersphere (T-hp), which follows the density:
$p_{X}(x^{(i)})\propto\frac{2}{\kappa(1-\mu^{T}x^{(i)})},$ (3)
@@
$N_{T}(\kappa,d)=\int_{\mathbb{S}^{d-1}}{\frac{2}{1-\kappa\mu^{T}x^{(i)}}}\mathrm{d}x=\frac{1}{\kappa}2^{\alpha+\beta-1}\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)},$ (4)
where $\alpha=\frac{d-1}{2}$, $\beta=\frac{d-3}{2}$, and $\Gamma(\cdot)$ represents the gamma function. Combined with Eq. 9 in Appendix A.2.1, the normalizer $N_{X}(d)$ of density $p_{X}(x;\mu)$ is:
$N_{X}(\kappa,d)=N_{T}(\kappa,d)\cdot A_{d-2}=\frac{2^{\alpha+\beta}\pi^{\beta}}{\kappa}\frac{\Gamma(\alpha)}{\Gamma(\alpha+\beta)}.$ (5)
@@
$p(x^{(i)})=N_{X}(d)^{-1}\frac{2}{\kappa(1-\mu^{T}x^{(i)})},\quad x^{(i)}\sim\text{Thp}(\mu).$ (6)
The detailed derivation process is provided in Appendix A.2.
@@
In terms of adapting to tailed concept drift, the Thp metric with a large concentration exhibits a light-tailed property, wherein the probability density function exhibits a faster rate of decay as the values increase, relative to the vMF metric, as illustrated in Figure 3. The high kurtosis of Thp is characterized that it yields high confidence only when the feature vector is sufficiently close to the center of the category, thereby minimizing the influence of head category samples on the tail category centers. Formally, $L_{\text{thp}}(\mu,x^{(i)})=\frac{2}{\kappa(1-\mu^{T}x^{(i)})+\epsilon}$ denotes the T-distributed metric on hypersphere, where $\epsilon$ is a non-zero value setting to $1$ to avoid the denominator of 0, and $L_{\text{vmf}}(\mu,x)=\exp(\kappa\mu^{T}{x})$ represents the vMF metric. Given an unit feature vector $x_{\text{head}}$ from the head categories, the gradient of the metric $L$ over the tailed category center $\mu_{\text{tail}}$ is $\frac{\partial L(\mu_{\text{tail}},x_{\text{head}})}{\partial\mu_{\text{tail}}}$. Due to $\mu^{T}_{\text{tail}}x_{\text{head}}\in[-1,1]$, when $\kappa\geqslant 1$, it is readily obtain that $\frac{\partial L_{\text{thp}}}{\partial\mu_{\text{tail}}}<\frac{\partial L_{\text{vmf}}}{\partial\mu_{\text{tail}}}$. Consequently, the light-tailed Thp distribution effectively counteracts the squeezing of tail categories caused by an overwhelming number of head samples, thereby alleviating the bias induced by the tailed concept drift.
Likewise, the Thp metric is directly applied to detect the OOD concept drift. A sample with a unit feature vector $x$ is deemed out-of-distribution if it lies at a relatively large distance from the in-distribution (ID) data in the eigenspace. Following Sun et al. (2022), a simple non-parametric KNN is adopted to partition the data into two sets (ID vs. OOD), which does not impose any distributional assumption on the feature space. Here the distance is the Thp metric with respect to the k-th nearest neighbor.